OpenStudy (anonymous):

Can someone please tell me if I got the right answer? x^2 + y^2= 25 x^2 -y^2 --- --- = 1 5^2 5^2

OpenStudy (kenljw):

x^2 + y^2 = r^2 r = 25 to find x, y you either need an angle, or given x or y

OpenStudy (kenljw):

r =5

OpenStudy (jdoe0001):

you got the right answer there, I said it hehe anyhow, yes, but is +, no - for the fractions, which btw are just a circle, or a "round ellipse" for that matter